Mrs. Kirby, 58, is a moderately active woman. She works in the garden almost daily, occasionally walks with her friends, and babysits her three very active young grandchildren every weekend. She eats a fairly standard diet every day: she has cereal with milk and coffee for breakfast, some kind of salad for lunch, and a piece of meat or fish with a healthy helping of veggies for dinner She usually skips desserts or has just a couple of strawberries. When asked to look at the food labels and calculate her intake of carbohydrates, fats, and proteins in grams over the entire day, on an average day, she came up with the following numbers: Carbohydrates: 243 g Fats: 41 g Proteins: 63 g Calculate what percentage of Mrs. Kirbys total energy intake came from carbohydrates. Round the percentage to the nearest whole number. View Available Hint(s) Submit Part B Energy Contribution of Fats After Mrs. Kirby received nutritional advice, her neighbor asked for her help with calculating the contribution of energy from fats in her diet For a long time, her doctor has been telling her to lower the fat intake below 30% of her total calories, but she does not really know how to do it-much less how to calculate it Mrs. Kirby explained how she did her calculations, and the neighbor showed her her numbers. Together they came up with total calorie intake of 2200 kcal and 63 g of fat What is the percentage contribution of energy from fat in Mrs. Kirbys neighbors diet? Round the percentage to the nearest whole number
media%2Ffe2%2Ffe2e8ec1-dcd2-42a6-8cfe-6f
Answer #1

A. Total energy uptake is the total number of calories consumed per day. It is the sum of the calories from protein, fats and carbohydrates.

(* denotes multiplication)

1 gm of protein has 4 Kcal energy

1 gm of carbohydrate has 4 Kcal energy

1 gm of fat has 9 Kcal energy different components is:

Energy obtained fro

Carbohydrates= 243 g * 4= 972 Kcal

Protein= 63 g *4= 252 Kcal

Fats= 41 g * 9= 369 Kcal

Total energy uptake= 972 +252 + 369= 1593 Kcal

Percentage of total energy intake from carbohydrates= Energy from carbohydrates/Total energy uptake *100= 972/1593= 61.01= 61%

Percentage of total energy intake from carbohydrates= 61%

B. Total energy uptake of Mrs Kirby’s friend is 2200 Kcal

Fat consumed= 63 g

Energy from fat= 63 g * 9= 567 Kcal

Percentage of total energy intake from fats= Energy from fats/Total energy uptake *100= 567/ 2200= 25.77= 26%

Percentage of total energy intake from fats for Mrs Kirby’s neighbour= 26%

Acceptable Macronutrient Distribution Range or AMDR, indicates the amount of carbohydrates, fats and proteins that is allowed for a person per day. The AMDR for fats is 20-30% of the total intake for women between 19-above 70 years of age. As the percentage of total energy intake from fats for Mrs Kirby’s neighbor is 26%, it is within the AMDR range as per her calorie intake.

Right answer is: Yes.